the third term in the expansion of (1/x+xlog10x)5 is 1 then x =
 

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
3rd term = 5C2(1/x)^3 (x log x)^2 = 1
 
(log x)^2 = x/10
 
log x =sqrt( x/10)
 
Hope it helps
Now Please check as is this one or